A capacitor of capacitance 3.0 uF is subjected to 2000 V potential difference across it's terminals. Calculate the energy stored in the capacitor
1
Expert's answer
2021-08-16T16:36:12-0400

Gives

"C=3\mu F"

V=2000V

We know that

"U=\frac{1}{2}CV^2"

Put value

"U=\frac{1}{2}\times3.0\times10^{-6}\times(2000)^2"


"U=\frac{1}{2}\times3.0\times10^{-6}\times4\times10^6=6J"